Suppose you wish to make a solenoid whose self-inductance is 1.2 mH. The inductor is to have a cross-sectional area of 2.00 10-3 m2 and a length of 0.053 m. How many turns of wire are needed?
Answer #1

Giver, Self Inductae = 102mH -3.2 Cross sec Area = 2X10 M Length = 0.053 m turns = ? we know L=NNA L = Inductance = 0.0012 H

IM = 2716 N=407107 put values, we get D.0012 - I XD XNX2X10 O. 053 0.0012 = 4 * 3.14 X 10 xnx2x103 0.053 - 6 pol2 = 12.5% 12

N = 2.53 * 104

N = 159

Number of turns would be 159

hope this helps , regards